OpenStudy (anonymous):

Find a vector equation of the plane that contains the origin and the point (2, -3, 2) and is perpendicular to the plane x+2y-z+3=0?

OpenStudy (anonymous):

\[\begin{matrix} \rightarrow \\ r\end{matrix}=s(1, 2, -1)+t(2, -3, 2)\]

OpenStudy (experimentx):

solve this equation \[ 2 - 3a + 2b = 0\\ 1 + 2a -c = 0\] and your plane is \[ x + ay + bz = 0\]